# Active Learning for Control-Oriented Identification of Nonlinear Systems

### v1

@@ 3 Proposed Algorithm and Main Result
$\frac{1}{\lambda_{\min}(\mathcal{H}(\phi^{\star}))},\left\|\mathcal{H}(\phi^{\star})\right\|,\sigma_{w},\log N,\log\frac{1}{\delta},\log B_{f},C_{\mathsf{Loja}}\bigg{)}.$
@@
The above result characterizes the excess control cost in terms of three key quantities. First, the term $\mathcal{H}(\phi^{\star})$ is the model-task Hessian, which describes how error in identification of the dynamics model parameters translates to the control cost. Second, is the inverse Fisher Information of the optimal exploration policy term, which measures a signal-to-noise ratio quantifying the hardness of parameter identification. Finally, the number of exploration episodes $N$ on the denominator captures the rate of decay from increasing the experimental budget.
@@
We note that in the setting where the dynamics model has linear dependence on the parameters, Wagenmaker et al. (2023) present a lower bound on the excess control cost achieved by any learner following the model based interaction protocol described in Section 2. The upper bound of Theorem 3.2 matches this lower bound up to universal constants, and the log term $\log\left(d_{\phi}/\delta\right)$. Future work will pursue general lower bounds that hold for dynamics models with a nonlinear dependence on the unknown parameter. We conjecture that in general, the upper bounds of Theorem 3.2 are tight up to universal constants and the term $\log\left(d_{\phi}/\delta\right)$.
We note that the burn-in time is currently polynomial in the relevant system parameters; however, we do not pursue tight burn-in times in this work. It may be possible to improve the dependence of the burn-in on various system quantities, e.g. by leveraging stability or reachability to obtain optimal dependence of the burn-in on $T$.
@@ 4 Proof Sketch
Full proof details may be found in the appendix. Here, we present a sketch. Our main result proceeds by demonstrating the following sub-steps. In these sub-steps, let $C$ be a polynomial of the problem parameters, as in the burn-in requirement of Theorem 3.2.
@@
    The coarse parameter estimation error decays gracefully with the total amount of data:
@@
    $\bigg{|}\!\operatorname{\mathrm{tr}}\left(\!\mathcal{H}(\hat{\phi}^{-})\Lambda_{\pi}^{-1}\!\right)\!-\!\operatorname{\mathrm{tr}}(\mathcal{H}(\phi^{\star})\mathbf{FI}^{\pi}(\phi^{\star})^{-1})\bigg{|}\!\leq\!C\!\left\|\hat{\phi}^{-}\!\!-\!\phi^{\star}\right\|,$ (19)
    with $\Lambda_{\pi}$ as in (15).
@@
    The call to the $\texttt{DOED}_{+}$ algorithm in 8 of Algorithm 1 returns a collection of policies whose mixture approximately minimizes the exploration objective (14):
    $\operatorname{\mathrm{tr}}\left(\mathcal{H}(\hat{\phi}^{-})\Lambda_{\pi_{\mathsf{mix}}}^{-1}\right)\leq\frac{C}{N^{\alpha}}\!+\!\min_{\tilde{\pi}\in\Pi_{\exp}}\operatorname{\mathrm{tr}}\Bigg{(}\mathcal{H}(\hat{\phi}^{-})\Lambda_{\tilde{\pi}}^{-1}\Bigg{)},$ (20)
    for $\Lambda_{\pi}$ as in (15). Demonstrating this result involves specializing the dynamic optimal exploration algorithm and its proof of correctness to our setting from Wagenmaker et al. (2023).
  4. 4.
@@
Using the above results, our argument proceeds according to the following series of inequalities applied to the excess cost:
@@
$\leq 40\log\left(\frac{24d_{\phi}}{\delta}\right)\frac{\operatorname{\mathrm{tr}}(\mathcal{H}(\phi^{\star})\mathbf{FI}^{\pi_{\mathsf{mix}}}(\phi^{\star})^{-1})}{N}+\frac{C}{N^{3/2}},$
where the first inequality follows by Footnote 3, and the second inequality follows by applying Theorem 3.1 with $H=\mathcal{H}(\phi^{\star})$ for the first term, and $H=I$ for the second term [^7] The condition that $\Sigma^{\pi_{\mathsf{mix}}}\succ 0$ in Theorem 3.1 can be shown from the success of DOED in (20)). . Next, it follows that
$\operatorname{\mathrm{tr}}(\mathcal{H}(\phi^{\star})\mathbf{FI}^{\pi_{\mathsf{mix}}}(\phi^{\star})^{-1})$ $\overset{(i)}{\leq}\frac{C}{N^{\alpha}}+\operatorname{\mathrm{tr}}\left(H(\hat{\phi}^{-})\operatorname{\mathbb{E}}_{\pi_{\mathsf{mix}}}^{\phi^{\star}}\left[(Df(x_{t},u_{t};\hat{\phi}^{-}))^{\top}Df(x_{t},u_{t};\hat{\phi}^{-})\right]^{-1}\right)$
$\overset{(ii)}{\leq}2\frac{C}{N^{\alpha}}+\min_{\tilde{\pi}\in\Pi_{\exp}}\operatorname{\mathrm{tr}}\Bigg{(}H(\hat{\phi}^{-})\operatorname{\mathbb{E}}_{\tilde{\pi}}^{\phi^{\star}}\left[(Df(x_{t},u_{t};\hat{\phi}^{-}))^{\top}Df(x_{t},u_{t};\hat{\phi}^{-})\right]^{-1}\Bigg{)}$
$\overset{(iii)}{\leq}3\frac{C}{N^{\alpha}}+\min_{\tilde{\pi}\in\Pi_{\exp}}\operatorname{\mathrm{tr}}\Bigg{(}H(\phi^{\star})\mathbf{FI}^{\tilde{\pi}}(\phi^{\star})^{-1}\Bigg{)},$
where inequality $(i)$ follows from (19) and (18), inequality $(ii)$ follows from (20), and inequality $(iii)$ follows from (19) and (18). The main result then follows by substituting the above bound into (21), and taking $N$ to exceed a polynomial burn-in time so the higher order terms become negligible.
@@ 5 Numerical Validation
We deploy ALCOI on an illustrative example to illustrate the benefits of active control-oriented exploration. For more experiments, and further details, see Appendix E. Consider the two dimensional system
@@ ACKNOWLEDGMENT
Ingvar Ziemann is supported by a Swedish Research Council international postdoc grant. George J. Pappas is supported in part by NSF Award SLES-2331880. Nikolai Matni is supported in part by NSF Award SLES-2331880 and NSF CAREER award ECCS-2045834.
@@ A.1 Consistency of Least Squares Parameter Estimation
Recent developments Ziemann and Tu (2022); Ziemann et al. (2023) have provided finite sample bounds for the $L_{2}$ prediction error of nonlinear least squares estimators in the presence of time dependent data with rates that do not depend on the mixing time of the data. Our first estimation bound instantiates these results in our setting under the assumption that the data collection policy is $(C_{\mathsf{Loja}},\alpha)$-Lojasiewicz. To express these bounds we define $\mathcal{G}=\mathopen{}\left\{f(\cdot,\cdot,\phi):\phi\in\mathbb{R}^{d_{\phi}}\right\}\mathclose{}$ and $\mathcal{G}_{\star}=\mathcal{G}-\mathopen{}\left\{f(\cdot,\cdot,\phi^{\star})\right\}\mathclose{}$. Using the Lipschitz bound on $f$, we find that the size $\mathcal{N}(\mathcal{G}_{\star},\left\|\cdot\right\|_{\infty},\varepsilon)$ of the the minimal cardinality $\varepsilon$-net covering $\mathcal{G}_{\star}$ is bounded as
$\log\mathcal{N}(\mathcal{G}_{\star},\left\|\cdot\right\|_{\infty},\varepsilon)\leq d_{\phi}\log\left(\frac{2L_{f}}{\varepsilon}+1\right).$ (22)
@@
$\tau_{\mathbf{Err}}(\delta)=\max\bigg{\{}\left(\frac{256\sigma_{w}^{2}}{T}\left(d_{\mathsf{X}}+d_{\phi}\log\left(2L_{f}TK+1\right)+\log\frac{1}{\delta}\right)\right)^{\frac{1}{4\alpha-1}},$ (23)
$\Bigg{(}4(L_{f}C_{\mathsf{Loja}}\alpha^{-1})^{4}\left(\log\left(4BK^{2\alpha}\right)+d_{\phi}\log\left(2L_{f}K^{2\alpha}+1\right)\right)+\log\frac{1}{\delta}\Bigg{)}^{\frac{1}{1-8\alpha+16\alpha^{2}}}\bigg{\}}.$
@@
$\mathbf{Err}_{\pi,\phi^{\star}}(\hat{\phi})\leq\frac{512\sigma_{w}^{2}}{TK}\left(d_{\mathsf{X}}+d_{\phi}\log\left(\frac{3L_{f}TK}{\delta}\right)\right).$
To prove this result, we first state several preliminary results. In order to do so, we introduce additional notation. Let $\mathcal{G}_{\star}^{s}=\mathopen{}\left\{bg:b\in[0,1],g\in\mathcal{G}_{\star}\right\}\mathclose{}$ be the star-shaped hull of $\mathcal{G}_{\star}$. [^8] A function class $\mathcal{G}$ is star-shaped if for all $b\in[0,1]$, $g\in\mathcal{G}\implies bg\in\mathcal{G}$. Since all elements of $\mathcal{G}_{\star}^{s}$ satisfy $\left\|f(\cdot,\cdot,\phi)\right\|_{\infty}\leq B$ for all $\phi\in\mathbb{R}^{d_{\phi}}$, we have from Lemma 4.5 of Mendelson (2002) that the size $\mathcal{N}(\mathcal{G}_{\star}^{s},\left\|\cdot\right\|_{\infty},\varepsilon)$ of a minimum cardinality $\varepsilon$-net of $\mathcal{G}_{\star}^{s}$ in the $\left\|\cdot\right\|_{\infty}$-norm satisfies
@@
Lemma A.4 (Lower isometry).
@@
$\operatorname{\mathbb{P}}\left[\sup_{g\in\mathcal{G}_{\star}^{s}\setminus B(r)}\frac{1}{TK}\sum_{t,k=1}^{T,K}\left\|g(X_{t}^{k},U_{t}^{k})\right\|^{2}-\frac{1}{8}\operatorname{\mathbb{E}}\left[\frac{1}{T}\left\|g(X_{t}^{1},U_{t}^{1})\right\|^{2}\right]\leq 0\right]\leq{\left|\mathcal{G}_{r}\right|}\exp\left(\frac{-Kr^{4-2a}}{4C_{\mathsf{hc}}}\right).$
@@
We may bound $\log{\left|\mathcal{G}_{r}\right|}$ using (24) and (22) as $\log{\left|\mathcal{G}_{r}\right|}\leq\log\left(\frac{4B}{r}\right)+d_{\mathsf{X}}\log\left(\frac{2L_{f}}{r}+1\right).$ Then choosing $r^{2}=\frac{1}{K^{a}}$, we find that by the requirement that
$K\geq\left(4C_{\mathsf{hc}}\left(\log\left(4BK^{a/2}\right)+d_{\phi}\log\left(2L_{f}K^{a/2}+1\right)\right)+\log\frac{1}{\delta}\right)^{\frac{1}{1-2a+a^{2}}},$
@@
$\sup_{g\in\mathcal{G}_{\gamma}}\frac{1}{TK}\left(\sum_{t,k=1}^{T,K}\left(4\langle g(X_{t}^{n},U_{t}^{n}),W_{t}^{k}\rangle-\frac{1}{2}\left\|g(X_{t}^{n},U_{t}^{n})\right\|^{2}\right)\right)$ $\leq 16\sigma_{w}^{2}\log\left(\frac{{\left|\mathcal{G}_{\gamma}\right|}}{\delta}\right)\leq 16\sigma_{w}^{2}\left(d_{\phi}\log\left(\frac{2L_{f}}{\gamma}+1\right)+\log\frac{1}{\delta}\right),$ (29)
where the final inequality followed from (22). Setting $\gamma=\frac{\sigma_{w}}{TK}$, we may combine (28) with (29) to conclude
$\sup_{g\in\mathcal{G}_{\star}}M_{T,K}(g)\leq\frac{12\sigma_{w}^{2}d_{\mathsf{X}}}{TK}+\frac{\sigma_{w}^{2}}{(TK)^{2}}+16\sigma_{w}^{2}\left(d_{\phi}\log\left(2L_{f}TK+1\right)+\log\frac{1}{\delta}\right)\leq\frac{32\sigma_{w}^{2}}{TK}\left(d_{\mathsf{X}}+d_{\phi}\log\left(2L_{f}TK+1\right)+\log\frac{1}{\delta}\right).$
@@
$K\geq\left(\frac{256\sigma_{w}^{2}}{T}\left(d_{\mathsf{X}}+d_{\phi}\log\left(2L_{f}TK+1\right)+\log\frac{1}{\delta}\right)\right)^{\frac{1}{a-1}},$
we may combing the above inequality with (26) and (27) to achieve the desired result. ∎
With this result in hand, we may immediately determine a high probability bound on the parameter estimation error as it appears in the bound in Footnote 3.
@@
$\left\|\hat{\phi}\!-\!\phi^{\star}\right\|^{2}\!\leq\!\left(\frac{512\sigma_{w}^{2}}{TN}\left(d_{\mathsf{X}}\!+\!d_{\phi}\log\left(\frac{3L_{f}TN}{\delta}\right)\right)\right)^{2\alpha}.$

### v2

@@ 3 Proposed Algorithm and Main Result
where $\xi=\beta\left(3+16\left(\frac{128d_{\phi}\left(\left\|\mathcal{H}(\phi^{\star})\right\|+\nu\right)}{\mu\left(\lambda_{\min}(\mathcal{H}(\phi^{\star}))+\nu\right)}+d_{\phi}\right)\right)$ as long as
$N\!$ $\geq\!\mathsf{poly}_{\alpha}\bigg{(}\!T,L_{f},L_{\mathsf{cost}},L_{\pi^{\star}},d_{\phi},d_{\mathsf{X}},\frac{1}{\mu},r_{\mathsf{cost}}(\phi^{\star}),r_{\theta}(\phi^{\star})$
$\frac{1}{\lambda_{\min}(\mathcal{H}(\phi^{\star}))+\nu},\left\|\mathcal{H}(\phi^{\star})\right\|,\nu,\sigma_{w},\sigma_{w}^{-1},\log N,$
$\log\frac{1}{\delta},\log B,C_{\mathsf{Loja}},\frac{1}{\gamma},\frac{1}{\beta}\bigg{)}.$
@@
The above result characterizes the excess control cost in terms of three key quantities. First, the term $\mathcal{H}(\phi^{\star})$ is the model-task Hessian, which describes how error in identification of the dynamics model parameters translates to the control cost. Second is the inverse Fisher information of the optimal exploration policy term, which measures a signal-to-noise ratio quantifying the hardness of parameter identification. Finally, the number of exploration episodes $N$ on the denominator captures the rate of decay from increasing the experimental budget.
@@
In the setting where the dynamics model has linear dependence on the parameters, Wagenmaker et al. (2023) present a lower bound on the excess control cost achieved by any learner following the model based interaction protocol described in Section 2. If we choose the free parameter in the upper bound as $\beta\leq\left(\frac{128d_{\phi}\left(\left\|\mathcal{H}(\phi^{\star})\right\|+\nu\right)}{\mu\left(\lambda_{\min}(\mathcal{H}(\phi^{\star}))+\nu\right)}+d_{\phi}\right)^{-1}$, then the upper bound of Theorem 3.2 matches this lower bound up to universal constants, and the term $\log\left(4/\delta\right)$. Future work will pursue general lower bounds that hold for dynamics models with a nonlinear dependence on the unknown parameter.
The burn-in time is currently polynomial in the relevant system parameters; however, we do not pursue tight burn-in times in this work. It may be possible to improve the dependence of the burn-in on various system quantities, e.g. by leveraging stability or reachability to obtain optimal dependence of the burn-in on $T$. We additionally draw attention to the utility of the parameter $\beta$ and the algorithm hyperparameters $\nu$ and $\gamma$ for navigating the tradeoff between a good burn-in time, and optimal rates. One can take $\beta$, $\nu$, $\gamma$ and $\varepsilon$ arbitrarily close to zero, meaning that the coefficient characterizing the excess cost can become arbitrarily close to $2+4\log\frac{8}{\delta}$. The cost of doing so is an increase in the burn-in time. A notable exception is the situation where $\mathcal{H}(\phi^{\star})\succ 0$, i.e. the setting where all parameters are necessary for control. In this case, one can take $\nu=0$, while the burn-in time must exceed a polynomial in $\frac{1}{\lambda_{\min}(\mathcal{H}(\phi^{\star}))}$.
@@ 4 Proof Sketch
Full proof details may be found in the appendix. Here, we present a sketch. Our main result proceeds by demonstrating the following sub-steps. In these sub-steps, let $C$ be a polynomial of the problem parameters and log of the reciprocal of the failure probability, $\delta$, as in the burn-in requirement of Theorem 3.2.
@@
    With high probability, the coarse parameter estimation error decays gracefully with the total amount of data:
@@
    $\bigg{|}\!\operatorname{\mathrm{tr}}\left(\!\mathcal{H}(\hat{\phi}^{-})\mathbf{FI}^{\pi}(\hat{\phi}^{-})^{-1}\!\right)\!-\!\operatorname{\mathrm{tr}}(\mathcal{H}(\phi^{\star})\mathbf{FI}^{\pi}(\phi^{\star})^{-1})\bigg{|}\!$ (15)
    $\leq\!C\!\left\|\hat{\phi}^{-}\!\!-\!\phi^{\star}\right\|.$
@@
Using the above results, our argument proceeds according to the following series of inequalities applied to the excess cost. With high probability,
@@
$\leq\xi(\delta)\frac{\operatorname{\mathrm{tr}}(\left(\mathcal{H}(\phi^{\star})+\nu I\right)\mathbf{FI}^{\pi_{\mathsf{mix}}}(\phi^{\star})^{-1})}{N}+\frac{C}{N^{3/2}},$
where the first inequality follows by Lemma 3.1, and the second inequality follows by applying Theorem 3.1 with $H=\mathcal{H}(\phi^{\star})+\nu I$ for the first term, and $H=I$ for the second term. The quantity $\xi(\delta)$ is a trades off the burn-in time and the final bound. In our analysis, it can become as small as $2+4\log\frac{1}{\delta}$. Next, it follows that
$\operatorname{\mathrm{tr}}(\left(\mathcal{H}(\phi^{\star})+\nu I\right)\mathbf{FI}^{\pi_{\mathsf{mix}}}(\phi^{\star})^{-1})$
$\overset{(i)}{\leq}\frac{C}{N^{\alpha}}+\operatorname{\mathrm{tr}}\left(\left(H(\hat{\phi}^{-})+\nu I\right)\mathbf{FI}^{\pi_{\mathsf{mix}}}(\hat{\phi}^{-}){-1}\right)$
$\overset{(ii)}{\leq}\!2\frac{C}{N^{\alpha}}\!+\!\frac{1}{1-\gamma}\inf_{\tilde{\pi}\in\Pi_{\exp}}\!\operatorname{\mathrm{tr}}\Bigg{(}\!\left(H(\hat{\phi}^{-})\!+\!\nu I\right)\mathbf{FI}^{\tilde{\pi}}(\hat{\phi}^{-})^{-1}\!\Bigg{)}$
$\overset{(iii)}{\leq}\!3\frac{C}{N^{\alpha}}\!+\!\frac{1}{1\!-\!\gamma}\!\inf_{\tilde{\pi}\in\Pi_{\exp}}\!\operatorname{\mathrm{tr}}\Bigg{(}\left(H(\phi^{\star})\!+\!\nu I\right)\mathbf{FI}^{\tilde{\pi}}(\phi^{\star})^{-1}\Bigg{)},$
where inequality $(i)$ follows from (15) and (14), inequality $(ii)$ follows from the definition of the policy $\pi_{\mathsf{mix}}$ and inequality $(iii)$ follows from (15) and (14). The main result then follows by substituting the above bound into (16), and taking $N$ to exceed a polynomial burn-in time so the higher order terms become negligible.
@@ 5 Numerical Validation
We deploy ALCOI on an illustrative example to illustrate the benefits of active control-oriented exploration. For more experiments, and further details, see Appendix D. Consider the two dimensional system
@@ ACKNOWLEDGMENT
BL and NM are supported by NSF Award SLES-2331880, NSF CAREER award ECCS-2045834 and AFOSR Award FA9550-24-1-0102. IZ is supported by a Swedish Research Council international postdoc grant. GP is supported in part by NSF Award SLES-2331880.
@@ A.1 Consistency of Least Squares Parameter Estimation
Recent developments Ziemann and Tu (2022); Ziemann et al. (2023) have provided finite sample bounds for the $L_{2}$ prediction error of nonlinear least squares estimators in the presence of time dependent data with rates that do not depend on the mixing time of the data. Our first estimation bound instantiates these results in our setting under the assumption that the data collection policy is $(C_{\mathsf{Loja}},\alpha)$-Lojasiewicz. To express these bounds we define $\mathcal{G}=\mathopen{}\left\{f(\cdot,\cdot,\phi):\phi\in\mathbb{R}^{d_{\phi}},\left\|\phi\right\|\leq B\right\}\mathclose{}$ and $\mathcal{G}_{\star}=\mathcal{G}-\mathopen{}\left\{f(\cdot,\cdot,\phi^{\star})\right\}\mathclose{}$. Using the Lipschitz bound on $f$, we find that the size $\mathcal{N}(\mathcal{G}_{\star},\left\|\cdot\right\|_{\infty},\varepsilon)$ of the the minimal cardinality $\varepsilon$-net covering $\mathcal{G}_{\star}$ is bounded as
$\log\mathcal{N}(\mathcal{G}_{\star},\left\|\cdot\right\|_{\infty},\varepsilon)\leq d_{\phi}\log\left(\frac{3BL_{f}}{\varepsilon}+1\right).$ (17)
@@
$\tau_{\mathbf{Err}}(\delta)=\max\bigg{\{}\left(\frac{256\sigma_{w}^{2}}{T}\left(d_{\mathsf{X}}+d_{\phi}\log\left(4BL_{f}TK/\sigma_{w}+1\right)+\log\frac{1}{\delta}\right)\right)^{\frac{1}{4\alpha-1}},$ (18)
$\Bigg{(}8(L_{f}C_{\mathsf{Loja}}\alpha^{-1})^{4}\left(d_{\phi}\log\left(4BL_{f}K^{2\alpha}+1\right)\right)+\log\frac{1}{\delta}\Bigg{)}^{\frac{1}{1-8\alpha+16\alpha^{2}}}\bigg{\}}.$
@@
$\mathbf{Err}_{\pi,\phi^{\star}}(\hat{\phi})\leq\frac{512\sigma_{w}^{2}}{TK}\left(d_{\mathsf{X}}+d_{\phi}\log\left(\frac{4L_{f}TK}{\sigma_{w}\delta}\right)\right).$
To prove this result, we first state several preliminary results. In order to do so, we introduce additional notation. Let $\mathcal{G}_{\star}^{s}=\mathopen{}\left\{bg:b\in[0,1],g\in\mathcal{G}_{\star}\right\}\mathclose{}$. This set is called the star-hull of $\mathcal{G}_{\star}$ [^5] The star-hull of $\mathcal{G}$ is the smallest set containing $\mathcal{G}$ that is star-shaped.. Since all elements of $\mathcal{G}_{\star}^{s}$ satisfy $\left\|f(\cdot,\cdot,\phi)\right\|_{\infty}\leq B$, we have from Lemma 4.5 of Mendelson (2002) that the size $\mathcal{N}(\mathcal{G}_{\star}^{s},\left\|\cdot\right\|_{\infty},\varepsilon)$ of a minimum cardinality $\varepsilon$-net of $\mathcal{G}_{\star}^{s}$ in the $\left\|\cdot\right\|_{\infty}$-norm satisfies
@@
Lemma A.4 (Lower isometry, Multi-trajectory Modification of Theorem 5.2 of Ziemann and Tu (2022)).
@@
$\operatorname{\mathbb{P}}\left[\forall g\in\mathcal{G}_{\star}^{s}\setminus B(r),\frac{1}{TK}\sum_{t,k=1}^{T,K}\left\|g(X_{t}^{k},U_{t}^{k})\right\|^{2}-\frac{1}{8}\operatorname{\mathbb{E}}\left[\frac{1}{T}\left\|g(X_{t}^{1},U_{t}^{1})\right\|^{2}\right]\leq 0\right]\leq{\left|\mathcal{G}_{r}\right|}\exp\left(\frac{-Kr^{4-2a}}{4C_{\mathsf{hc}}}\right).$
@@
We may bound $\log{\left|\mathcal{G}_{r}\right|}$ using (19) and (17) as $\log{\left|\mathcal{G}_{r}\right|}\leq\log\left(\frac{4B}{r}\right)+d_{\phi}\log\left(\frac{3BL_{f}}{r}+1\right).$ Then choosing $r^{2}=\frac{1}{K^{a}}$, we find that by the requirement that
$K\geq\left(8C_{\mathsf{hc}}\left(d_{\phi}\log\left(4L_{f}K^{a/2}\right)\right)+\log\frac{1}{\delta}\right)^{\frac{1}{1-2a+a^{2}}},$
@@
$\sup_{g\in\mathcal{G}_{\gamma}}\frac{1}{TK}\left(\sum_{t,k=1}^{T,K}\left(4\langle g(X_{t}^{n},U_{t}^{n}),W_{t}^{k}\rangle-\frac{1}{2}\left\|g(X_{t}^{n},U_{t}^{n})\right\|^{2}\right)\right)$ (24)
$\leq 16\sigma_{w}^{2}\log\left(\frac{{\left|\mathcal{G}_{\gamma}\right|}}{\delta}\right)\leq 16\sigma_{w}^{2}\left(d_{\phi}\log\left(\frac{3BL_{f}}{\gamma}+1\right)+\log\frac{1}{\delta}\right),$
where the final inequality followed from (17). Setting $\gamma=\frac{\sigma_{w}}{TK}$, we may combine (23) with (LABEL:eq:_finite_offset_complexity_bound) to conclude
$\sup_{g\in\mathcal{G}_{\star}}M_{T,K}(g)$ $\leq\frac{12\sigma_{w}^{2}d_{\mathsf{X}}}{TK}+\frac{\sigma_{w}^{2}}{(TK)^{2}}+16\sigma_{w}^{2}\left(d_{\phi}\log\left(3BL_{f}TK/\sigma_{w}+1\right)+\log\frac{1}{\delta}\right)$
$\leq\frac{32\sigma_{w}^{2}}{TK}\left(d_{\mathsf{X}}+d_{\phi}\log\left(3BL_{f}TK/\sigma_{w}+1\right)+\log\frac{1}{\delta}\right).$
@@
$K\geq\left(\frac{256\sigma_{w}^{2}}{T}\left(d_{\mathsf{X}}+d_{\phi}\log\left(4L_{f}TK/\sigma_{w}\right)+\log\frac{1}{\delta}\right)\right)^{\frac{1}{a-1}},$
we may combine the above inequality with (21) and (22) to achieve the desired result. ∎
With this result in hand, we may immediately determine a high probability bound on the parameter estimation error as it appears in the bound in Lemma 3.1.
@@
$\left\|\hat{\phi}\!-\!\phi^{\star}\right\|^{2}\!\leq\!\left(\frac{512\sigma_{w}^{2}}{TK}\left(d_{\mathsf{X}}\!+\!d_{\phi}\log\left(4BL_{f}TK\right)\right)+\log\frac{1}{\delta}\right)^{2\alpha}.$